# WeAudit: Scaffolding User Auditors and AI Practitioners in Auditing Generative AI

### v3

@@ 5.1. Study with User Auditors
To investigate how end users audit AI systems using WeAudit, we conducted a user study in which 45 user auditors used WeAudit over a three-week period. We held two synchronous study sessions, with separate groups of participants (N1=15, N2=30). In each session, participants first received a 10-minute onboarding presentation to introduce them to the concept of biases and harms in text-to-image GenAI, and to the WeAudit interface. Participants then used WeAudit to conduct audits individually for 40 minutes. Afterward, they spent 10 minutes completing a post-audit survey and sharing their experiences using WeAudit. To better understand auditors’ experience of using WeAudit, we then facilitated a 30-minute group discussion, during which the research team took notes on the feedback provided. Following this initial study session, auditors were given three weeks to voluntarily continue using WeAudit’s discussion forum in their spare time. Informed by prior work in crowdsourcing (Bernstein et al., 2010), to provide a preliminary assessment of users’ audit reports, we then enlisted other users to verify reports. The verification process is described in Section 4.2.7.
@@ 5.1.1. Participants
We conducted our user study with 45 students at a US-based university. The students spanned different academic backgrounds, including computer science, engineering, design, public policy, mechanical engineering, and social sciences. Aggregate statistics of the user auditor participants are provided in Table 4 under the Appendix A.1. For the follow-up survey, we reached out to 37 participants who indicated interest in being contacted for follow-up activities. 17 participants responded to the follow-up survey. We chose to run our initial study with university students in order to avoid burdening marginalized communities at this stage of the research (Dourish et al., 2020; Pierre et al., 2021). Rather, our goal in conducting this study was to develop an initial understanding of WeAudit’s strengths and limitations, to identify opportunities for improvement prior to deploying the system more broadly. This user study allowed us to present industry GenAI practitioners with real user audit outputs in our interviews, enabling them to evaluate not only the WeAudit workflow and system, but also the concrete audit outputs produced through WeAudit. In the Discussion section (see Section 7), we expand on how readers should interpret our results and outline future work needed to further evaluate WeAudit’s effectiveness with broader population of user auditors.
@@ 5.2. Interviews with Industry GenAI Practitioners
We next conducted semi-structured interviews with 10 industry GenAI practitioners who evaluate and audit GenAI systems as part of their work. In line with prior work (Wang et al., 2024; Deng et al., 2022; Kaur et al., 2020), our goal in these interviews was to solicit their reflections on the WeAudit workflow, grounded in the outcomes of our user study, including how they might envision using the workflow as part of their GenAI development. In each interview, we first asked participants to describe their current work on GenAI evaluation, and then provided a brief overview of the WeAudit workflow and system. We then asked participants to evaluate different components of the WeAudit system and to envision how they might adapt certain features of WeAudit or integrate the WeAudit workflow into their existing GenAI development pipeline.
@@ 6.3.2. Opportunities for improving the verification process
In our interviews with practitioners, they were able to focus on audit reports that had gone through a crowd verification step. For the majority of user audit reports, verifiers agreed that they could “understand why the reporter finds this AI behavior harmful based on their report,” with an average of 80.35% agreement ($\sigma$=1.64%) across all reports. For reports with low agreement across verifiers, we found that the most common reason given by crowd worker verifiers for disagreeing with a report was due to ambiguity (as indicated by low ’clarity’ ratings for these reports), and not necessarily substantive disagreements (Chen and Zhang, 2023; Kuo et al., 2024). To validate the verifiers’ results, the research team compared verifiers’ agreement percentages with the research team’s own ratings for each report (see Section 5.1). While we found strong overall alignment with the research team’s ratings, we also observed a small number of cases where all members of our team understood the harm that a report identified, yet a majority of verifiers agreed that they could not understand why the user found the AI output harmful. We hypothesize these may have been cases where the user auditor came from a marginalized group that was underrepresented among verifiers. When we presented these observations to practitioners, they raised concerns about “how to ensure the aggregation mechanism in crowdsourcing does not further marginalize minority voices” (P05). In the next section, we discuss how the discussion function can complement the verification process by providing additional insights.
@@ 6.4. Enhancing Understanding of Audit Findings through Collective Discussion
In the three-week discussion phase for the study with user auditors, following the initial auditing session, 17 users posted a total of 62 comments, across 43 audit reports, on the auditing discussion forum. Analyzing the content of these discussions, we identified four main types of comments, each serving a different function: (1) Expressing surprise, (2) Providing additional evidence on harms, (3) Providing counterpoints or disagreements, and (4) Providing potential solutions to mitigate harms. Table 2 provides examples of comments in each category.
@@ 6.4.2. Collective discussions can enrich sensemaking and prioritization of audit reports
In addition, when cross-referencing the verification results from crowdworkers with users’ discussion comments, multiple practitioners (P01, P05, P07, P10, and P11) noted that the disagreements surfaced in the discussion could complement the majority vote-based verification approaches that their teams currently used. Discussion comments helped practitioners better understand why people might disagree. For example, P07 suggested that “discussion can surface disagreements and allow them to provide their rationales through conversations, which is better than the crowdsourcing evaluation you just showed me, especially for those [that] have high disagreement.” This points to design opportunities to help practitioners efficiently draw actionable insights from potentially large numbers of user discussions in user-engaged auditing processes.
@@ 7. Study Limitations
As mentioned in Section 5.1, while our formative study was conducted with a broader population of users, we chose to run our initial user study with college students. While this study provides insights into the usability and usefulness of the WeAudit workflow, an important direction for future work is to explore how a broader population of user auditors would use the WeAudit workflow and what additional forms of support and scaffolding they might need. Prior research in HCI, psychology, and AI has demonstrated that factors such as cultural background, technical expertise, and prior AI experience can significantly impact individuals’ sensitivity to and assessment of harmful AI biases and discrimination (Kingsley et al., 2024; Mun et al., 2024). Therefore, future work should engage broader groups in evaluating the usefulness and usability of the WeAudit system. Engaging a broader range of user auditors will help us understand how these diverse factors may influence auditors’ approaches to user-engaged AI audits. Building on insights from our user study, in the next phase of our research we plan to evaluate an improved version of WeAudit to with a broader population.
@@ 8. Discussion
More than a decade ago, in CSCW 2013, Kittur et al. asked a provocative question in their position paper “The Future of Crowd Work”: “Can we envision a future of crowd work where we would like our children to participate?” (Kittur et al., 2013). Inspired by this, we pose a similar question “can we envision a future of user AI audit where we would like our children (and ourselves) to participate?” As argued by Metaxa et al., the nature of algorithm auditing, especially involving users, is rooted in activism (Metaxa et al., 2021). Past research shows that users often advocate for marginalized groups, expressing solidarity through their auditing work (DeVos et al., 2022). The “We” in WeAudit stands for individuals impacted by AI systems, particularly those from marginalized communities. Engaging users in auditing AI is only the beginning of the WeAudit vision; we ultimately aim to empower these users to hold AI companies accountable. How can we, as a community, design a responsible future of user-engaged AI audit and user-empowered accountability?
In this section, we first discuss ways to ethically incentivize and sustain diverse participation in user AI auditing. We then outline how future work could cautiously incorporate AI technology into user AI auditing, based on our user study results. Finally, we highlight potential asymmetric power dynamics between user auditors and AI practitioners, calling for further research and policy development.
@@ 8.1. Incentivizing and Sustaining Participation in User-engaged AI Audits
As user-engaged AI audits become increasingly vital for identifying and addressing AI harms, motivating diverse and sustained participation remains a central challenge. Our study results indicate that the sense of community and identity that users feel within the platform plays a significant role in motivating user auditors (Section 6.2). Drawing on insights from CSCW research, initial recruitment for user-engaged AI audit can benefit from existing social groups (Kraut and Resnick, 2011; Burke et al., 2011; Lampe et al., 2010); in our study, which can naturally bolster ongoing engagement. To amplify this intrinsic motivations, future work could enhance the WeAudit forum by allowing user auditors to share specific cases on social media, inviting others to also participate in AI auditing. By incorporating social features that enable users to advocate for AI improvements within their personal networks, WeAudit could support users auditors to initiate collective advocacy. As seen in prior research on “user-driven algorithm audit,” public sharing of audit findings can drive tangible changes in AI systems (DeVos et al., 2022). Leveraging motivations such as “compassion” and the desire to improve society could further empower individuals to participate in the collective action of AI auditing (Kraut and Resnick, 2011).
However, intrinsic motivation shouldn’t be the only driver for users to engage in AI audits. User auditors, especially those from marginalized communities (Pierre et al., 2021; Li et al., 2023; Deng et al., 2024), should receive extrinsic rewards through fair compensation from companies that benefit from their labor. Following calls from prior research on crowd workers and content moderators (Hara et al., 2018; Kittur et al., 2013; Dosono and Semaan, 2019; Gray and Suri, 2019), companies should consider fair compensation for user auditors’ contributions, especially the “invisible labor” that goes into conducting audits (see Section 6.5). Policymakers can draw from content moderation research to design policies that encourage companies to establish protective mechanisms (Steiger et al., 2021), safeguarding user auditors from harmful exposure and fostering a safe, sustainable environment for user engagement in AI audits.
@@
Our findings indicate that some user auditors preferred the role of commenter over that of an auditor, choosing to provide feedback on others’ reports rather than conducting their own audits. (Section 6.4.1). This aligns with Li et al.’s findings that users often gravitate towards specific contributions in auditing–e.g. some were hypothesizing about AI harms, while others focused on amplifying or escalating audit outcomes (Li et al., 2023).
Indeed, enabling a broader range of participation levels—what is often referred to as a “low floor, high ceiling” design—can engage diverse users who may have varying time and interests to dedicate to different parts of auditing (Reynante et al., 2021). To this end, similar to the “ladder of citizen participation” model in civic engagement developed by Arnstein (Arnstein, 1969), a user-engaged AI audit system should also accommodate different levels and types of involvement. For instance, drawing from the concept of “microtask” in crowdsourcing research (Dow et al., 2012), WeAudit could introduce “microinteractions,” allowing users to select “sub-activities” within the WeAudit workflow according to their availabilities and interests. By further breaking down the workflow into smaller, modular tasks—such as generating prompts, inspecting images, issuing reports, and discussing or verifying others’ reports(cf. Claire et al., 2024)WeAudit can cater to users who might prefer (or only have availability for) brief, targeted interactions rather than long-form audits.
@@ 8.2. Designing Effective AI-in-the-Loop User-Engaged AI audits
In line with the “AIs Guiding Crowds” research agenda outlined by Kittur et al. (Kittur et al., 2013), recent research has increasingly explored the use of emerging generative AI technologies to assist developers in testing AI systems, brainstorming, and creating incident cases (Buçinca et al., 2023; Rastogi et al., 2023; Pang and Reinecke, 2023; Park et al., 2022; Kieslich et al., 2024; Pang et al., 2024; Wang et al., 2024; Zaydi and Maleh, 2024; Dev et al., 2023; Chiu et al., 2024; Radharapu et al., 2023). In the current WeAudit workflow, we opted not to include AI-powered features, given that prior studies have highlighted the potential limitations and risks associated with using generative AI for tasks like AI auditing, red-teaming, and impact assessment (Buçinca et al., 2023; Rastogi et al., 2023; Chiu et al., 2024; Radharapu et al., 2023). However, our user study results point to several promising directions for cautiously incorporating AI-powered features to enhance key activities in the user-engaged AI audit pipeline.
@@
As practitioners repeatedly emphasized when envisioning incorporating WeAudit, user-engaged audits should ideally be integrated directly into AI products and services to better contextualize the use case (Section 6). One challenge in this integration lies in determining the optimal timing and context for nudging users to provide audit feedback during regular tasks. To this end, future work could build on recent advancements in generative AI to develop personalized, context-aware task routing for user AI audits, drawing from prior CSCW research on “intelligent task routing” (Cosley et al., 2007). This would also benefit sustaining user engagement (see Section 8.1) by automatically providing users auditors tasks they are most interested in.
@@ Acknowledgements.
We thank all participating industry practitioners and user auditors for making this work possible.

### v4

@@ 5.1. Study with User Auditors
To investigate how end users audit AI systems using WeAudit, we conducted a user study in which 45 user auditors used WeAudit over a three-week period. Participants first received a 10-minute onboarding presentation to introduce them to the concept of biases and harms in text-to-image GenAI, and to the WeAudit interface. Participants then used WeAudit to conduct audits individually for 40 minutes. Similar to formative study (Section 3.1), we first provided participants with a diverse set of expert-led audits examples on T2I models. In addition, we specifically mentioned that whether the AI output is “harmful” can be subjective based on their own experiences and perceptions. Afterward, they spent 10 minutes completing a post-audit survey and sharing their experiences using WeAudit. To better understand auditors’ experience of using WeAudit, we then facilitated a 30-minute group discussion, during which the research team took notes on the feedback provided. Following this initial study session, auditors were given three weeks to voluntarily continue using WeAudit’s discussion forum in their spare time. Informed by prior work in crowdsourcing (Bernstein et al., 2010), to provide a preliminary assessment of users’ audit reports, we then enlisted other users to verify reports. The verification process is described in Section 4.2.7.
@@ 5.1.1. Participants
We conducted our user study with 45 students at a US-based university. We held two synchronous study sessions, with separate groups of participants (N1=15, N2=30), from two different classrooms which allowed us to observe interaction dynamics beyond a single group. The students spanned different academic backgrounds, including computer science, engineering, design, public policy, mechanical engineering, and social sciences. Aggregate statistics of the user auditor participants are provided in Table 3 under the Appendix A.1. For the follow-up survey, we reached out to 37 participants who indicated interest in being contacted for follow-up activities. 17 participants responded to the follow-up survey. We chose to run our initial study with university students in order to avoid burdening marginalized communities at this stage of the research (Dourish et al., 2020; Pierre et al., 2021). Rather, our goal in conducting this study was to develop an initial understanding of WeAudit’s strengths and limitations, to identify opportunities for improvement prior to deploying the system more broadly. This user study allowed us to present industry GenAI practitioners with real user audit outputs in our interviews, enabling them to evaluate not only the WeAudit workflow and system, but also the concrete audit outputs produced through WeAudit. In the Discussion section (see Section 7), we expand on how readers should interpret our results and outline future work needed to further evaluate WeAudit’s effectiveness with broader population of user auditors.
@@ 5.2. Interviews with Industry GenAI Practitioners
We next conducted semi-structured interviews with 10 industry GenAI practitioners who evaluate and audit GenAI systems as part of their work. Our goal in these interviews was to solicit their reflections on the WeAudit workflow, grounded in the outcomes of our user study, including how they might envision using the workflow as part of their GenAI development. In each interview, we first asked participants to describe their current work on GenAI evaluation, and then provided a brief overview of the WeAudit workflow and system. We then asked participants to evaluate different components of the WeAudit system and to envision how they might adapt certain features of WeAudit or integrate the WeAudit workflow into their existing GenAI development pipeline.
@@ 6.3.2. Opportunities for improving the verification process
In our interviews with practitioners, they were able to focus on audit reports that had gone through a crowd verification step. For the majority of user audit reports, verifiers agreed that they could “understand why the reporter finds this AI behavior harmful based on their report,” with an average of 80.35% agreement ($\sigma$=1.64%) across all reports. For reports with low agreement across verifiers, we found that the most common reason given by crowd worker verifiers for disagreeing with a report was due to ambiguity (as indicated by low “clarity” ratings for these reports), and not necessarily substantive disagreements (Chen and Zhang, 2023; Kuo et al., 2024). To validate the verifiers’ results, the research team compared verifiers’ agreement percentages with the research team’s own ratings for each report (see Section 5.1). While we found strong overall alignment with the research team’s ratings, we also observed a small number of cases where all members of our team understood the harm that a report identified, yet a majority of verifiers agreed that they could not understand why the user found the AI output harmful. We hypothesize these may have been cases where the user auditor came from a marginalized group that was underrepresented among verifiers. When we presented these observations to practitioners, they raised concerns about “how to ensure the aggregation mechanism in crowdsourcing does not further marginalize minority voices” (P05). In the next section, we discuss how the discussion function can complement the verification process by providing additional insights.
@@ 6.4. Enhancing Understanding of Audit Findings through Collective Discussion
In the three-week discussion phase for the study with user auditors, following the initial auditing session, 17 users posted a total of 62 comments, across 43 audit reports, on the auditing discussion forum.
Analyzing the content of these discussions, we identified four main types of comments, each serving a different function: (1) Expressing surprise, (2) Providing additional evidence on harms, (3) Providing counterpoints or disagreements, and (4) Providing potential solutions to mitigate harms. Table 1 provides examples of comments in each category.
@@ 6.4.2. Collective discussions can enrich sensemaking and prioritization of audit reports
In addition, when cross-referencing the verification results from crowdworkers with users’ discussion comments, multiple practitioners (P01, P05, P07, P10, and P11) noted that the disagreements surfaced in the discussion could complement the majority vote-based verification approaches that their teams currently used. For example, F14 provided several comments explaining why they disagreed with others about the AI outputs being “harmful,” along with their detailed rationales (See Table 1). These discussion comments helped practitioners better understand why people might disagree. For example, P07 suggested that “discussion can surface disagreements and allow them to provide their rationales through conversations, which is better than the crowdsourcing evaluation you just showed me, especially for those [that] have high disagreement.” This points to design opportunities to help practitioners efficiently draw actionable insights from potentially large numbers of user discussions in user-engaged auditing processes.
@@ 7. Study Limitations
As mentioned in Section 5.1, while our formative study was conducted with a broader population of users, we chose to run our initial user study with college students. While this study provides insights into the usability and usefulness of the WeAudit workflow, an important direction for future work is to explore how a broader population of user auditors would use the WeAudit workflow and what additional forms of support and scaffolding they might need. Prior research in HCI, psychology, and AI has demonstrated that factors such as cultural background, technical expertise, and prior AI experience can significantly impact individuals’ sensitivity to and assessment of harmful AI biases and discrimination (Kingsley et al., 2024; Mun et al., 2024). Therefore, future work should engage broader groups in evaluating the usefulness and usability of the WeAudit system. Engaging a broader range of user auditors will help us understand how these diverse factors may influence auditors’ approaches to user-engaged AI audits.
@@ 8. Discussion
More than a decade ago, in CSCW 2013, Kittur et al. asked a provocative question in their position paper “The Future of Crowd Work”: “Can we envision a future of crowd work where we would like our children to participate?” (Kittur et al., 2013). Inspired by this, we pose a similar question: “Can we envision a future of user-engaged AI auditing where we would like our children (and ourselves) to participate?” As argued by Metaxa et al., the nature of algorithm auditing, especially when involving users, is rooted in activism (Metaxa et al., 2021). Past research shows that users often advocate for marginalized groups, expressing solidarity through their auditing work (DeVos et al., 2022). The “We” in WeAudit stands for individuals impacted by AI systems, particularly those from marginalized communities. Engaging users in auditing AI is only the beginning of the WeAudit vision; we ultimately aim to empower these users to hold AI companies accountable. How can we, as a community, design a responsible future of user-engaged AI auditing?
In this section, we first discuss ways to ethically incentivize and sustain diverse participation in user-engaged AI auditing. We then outline how future work could cautiously incorporate AI technology into user-engaged AI auditing, based on our findings. Finally, we highlight potential asymmetric power dynamics between user auditors and AI practitioners, calling for further research and policy development.
@@ 8.1. Incentivizing and Sustaining Participation in User-Engaged AI Audits
Building on insights from our user study, in the next phase of our research, we plan to evaluate an improved version of WeAudit with a broader population. We also plan to launch WeAudit as a publicly available web platform for use by (1) AI practitioners and researchers interested in conducting user-engaged AI audits; (2) users interested in getting involved in user-engaged AI audits, either on a volunteer or paid basis; (3) researchers interested in conducting research on user-engaged AI auditing; and finally, (4) educators interested in using WeAudit as an educational resource for enhancing AI literacy regarding the societal impact of AI systems (e.g., (Shen et al., 2021a; Widder et al., 2024)), inspired by our finding that user auditors in our study reported increased awareness and understanding of AI harms (Section 6.6).
As user-engaged AI audits become increasingly vital for identifying and addressing AI harms, motivating diverse and sustained participation remains a central challenge. Our findings indicate that the sense of community and identity that users feel within the platform , often refereed as “intrinsic motivation” in CSCW research, plays a significant role in motivating user auditors (Section 6.2). Drawing on insights from prior CSCW work, initial recruitment for user-engaged AI audits can benefit from existing social groups to leverage established user bases that have shared interests or identities (Kraut and Resnick, 2011; Burke et al., 2011; Lampe et al., 2010). To amplify these intrinsic motivations, future work could enhance the WeAudit forum by allowing user auditors to share specific cases on social media, inviting others to also participate in AI auditing. By incorporating social features that enable users to advocate for AI improvements within their personal networks, WeAudit could help user auditors to collectively push for changes (DeVos et al., 2022).
However, intrinsic motivations should not be the only drivers for users to engage in AI auditing. User auditors, especially those from marginalized communities (Pierre et al., 2021; Li et al., 2023; Deng et al., 2024b), should receive extrinsic rewards through fair compensation from companies that benefit from their labor. Following calls from prior research on crowd workers and content moderators (Hara et al., 2018; Kittur et al., 2013; Dosono and Semaan, 2019; Gray and Suri, 2019), companies should consider fair compensation for user auditors’ contributions, especially the “invisible labor” that goes into conducting audits (see Section 6.5). Policymakers can draw from content moderation research to design policies that encourage companies to establish protective mechanisms (Steiger et al., 2021), safeguarding user auditors from harmful exposure and fostering a sustainable environment for user engagement in AI audits.
@@
Our findings indicate that some user auditors preferred the role of commenter over that of an auditor, choosing to provide feedback on others’ reports rather than conducting their own audits. (Section 6.4.1). Indeed, enabling a broader range of participation levels—what is often referred to as a “low floor, high ceiling” design—can engage diverse users who may have varying time and interests to dedicate to different parts of auditing (Reynante et al., 2021). To this end, a user-engaged AI audit system should also accommodate different levels and types of involvement. For example, drawing on the concept of “microtasks” in crowdsourcing research (Dow et al., 2012), WeAudit could allow users to select specific sub-activities within the WeAudit workflow according to their availability, interests, and specific domain or cultural expertise. By further breaking down the workflow into smaller, modular tasks—such as generating prompts, inspecting images, issuing reports, and discussing or verifying others’ reports—WeAudit can cater to users who might prefer (or only have availability for) brief, targeted interactions rather than long-form audits. This is particularly important when WeAudit and other future user auditing systems are deployed for broad public use, as past research shows that users often gravitate toward making specific kinds of contributions in auditing (Li et al., 2023).
@@ 8.2. Designing Effective AI-in-the-Loop User-Engaged AI audits
In line with the “AIs Guiding Crowds” research agenda outlined by Kittur et al. (Kittur et al., 2013), recent research has increasingly explored the use of emerging generative AI technologies to assist developers in testing AI systems, brainstorming, and creating incident cases (Buçinca et al., 2023; Rastogi et al., 2023; Pang and Reinecke, 2023; Park et al., 2022; Kieslich et al., 2024; Pang et al., 2024; Wang et al., 2024). In the current WeAudit workflow, we opted not to include AI-powered features, given that prior studies have highlighted the potential limitations and risks associated with using generative AI for tasks like AI auditing, red-teaming, and impact assessment (Buçinca et al., 2023; Rastogi et al., 2023; Zaydi and Maleh, 2024; Dev et al., 2023; Chiu et al., 2024; Radharapu et al., 2023). However, our findings point to several promising directions for thoughtfully incorporating AI-powered features to enhance key activities in the user-engaged AI audit pipeline.
@@
As practitioners repeatedly emphasized when envisioning incorporating WeAudit, user-engaged audits should ideally be integrated directly into AI products and services to better contextualize the use case (Section 6). One challenge in this integration lies in determining the optimal timing and context for nudging users to provide audit feedback during regular tasks. To this end, future work could build on recent advancements in generative AI to develop personalized, context-aware task routing for user-engaged AI audits, drawing from prior CSCW research on “intelligent task routing” (Cosley et al., 2007). This would also benefit sustaining user engagement (see Section 8.1) by automatically providing users auditors tasks they are most interested in.
@@ Acknowledgements.
We thank all participating industry practitioners and users for making this work possible.